# {Authlist}

### v1

@@ 3 Monte Carlo simulations
The Monte Carlo simulations were performed using three of the physics lists available in Geant4, as well as FLUKA (fluka 4-2.1). The model was simulated in Geant4.9.6.p04, in Geant4.10.05.p01, and the most recent one, in Geant4.11.2.1. The four models adopt different approaches when describing antiproton-nucleus annihilation at rest.
@@
For the hodoscope a straightforward threshold on the deposited energy was implemented to define a hit (0.7 MeV for the bars and of 0.55 MeV for fibres), determined from the deposited energy by simulated cosmic rays. Both Geant4 and FLUKA simulations consistently indicate that most of the tracks in the hodoscope are from charged pions. However, a significant portion is ascribed to other high–energy particles crossing the four layers of the hodoscope. In FTFP, CHIPS and INCL their contribution remains constant at 35%, 50% and 40% respectively, irrespective of the target. In contrast, for FLUKA the values range between 25% and 35%, increasing with the mass of the nucleus. In CHIPS and FTFP models 95% of the non-pion tracks are generated by electrons or positrons, each with tens of MeV of energy, resulting from the conversion of -rays emitted from the decay of from annihilation. In FLUKA however, the surplus of 3D reconstructed tracks is primarily attributed to high-energy protons. In INCL most of the non-pion tracks originate from electrons or positron, but proton contribution is significant with 40%. Fig. LABEL:fig:MIPs_distros_gold_sims shows the distribution of the number of 3D tracks per annihilation event for -Au, across the four models. Additionally, the graph includes a breakdown of the particles fully traversing the hodoscope, depositing in each layer energy above the defined threshold. It is evident that the distribution of charged pions alone does not reproduce the number of 3D reconstructed tracks, underscoring the contribution of other particles crossing the detector.
@@
The Timepix3 simulated data were digitised using the modular simulation framework, applying the same bias voltage and charge threshold as during the actual measurements [73, 74]. The clusters generated from MIPs in the cosmic rays simulations reproduced the morphology and the total cluster energy of the measured ones, as observed in Fig. LABEL:fig:cosmics_alphas_digia. On the other hand, the digitisation could not consistently reproduce the measured signal from HIPs, due to features of the Timepix3 ASIC, such as the halo, plasma and the volcano effect [58, 75, 76, 77, 78]. These effects arise from substantial energy deposits, and simulating them accurately is challenging because HIPs do not always trigger each of them. The halo (or "skirt"), indicating induced low energy signal in the neighboring pixels is consistently present. However, when heavier or more energetic ions deposit a significant amount of charge over short distances, two additional phenomena may occur: the plasma effect, which extends the charge collection time and consequently increases the track width, or the volcano effect, where pixels with energy deposits exceeding 500 keV register significantly lower values. The measured cluster energy is then notably lower than the actual deposited energy in the silicon sensor. In addition, the stochastic nature, in particular of the volcano effect, makes their qualitative and quantitative replication in the digitisation unfeasible. Therefore, these effects had to be individually addressed in the data analysis.
@@
The halo pixels, which are easily identifiable due to the very low energy deposits, were excluded implementing an energy cutoff of 5 keV, based on the pixel energy distribution derived from data. Fig. LABEL:fig:cosmics_alphas_digib shows a good agreement of the energy deposition by particles from a source between data (after halo removal) and Geant4 simulation, both before and after digitisation, despite the significantly larger clusters observed in data due to the presence of halo pixels. The small discrepancy is likely attributed to the uncertainties in measuring the distance between the source and the detector during the measurements.
@@ 4.1 Multiplicities
The charged pion multiplicity in antiproton-nucleus annihilation at rest is influenced by the final state interactions of the primary pions with the remaining nucleus. The number of pions that interact with the nucleus is predominantly determined by simple geometric factors, such as the proximity of the annihilation point to the nuclear surface and the size (mass) of the nucleus. The number of surviving, and detectable pions is related to the primary pions through:
@@
Due to the limited geometrical acceptance and efficiency of the hodoscope, only a fraction of the surviving charged pions are measured in this experiment. The distributions of MIP particles presented through the number of reconstructed 3D tracks in the hodoscope are shown in Fig. LABEL:fig:MIPs_distros. No model reproduces accurately the distributions obtained from the measurements for all three targets, with FLUKA, CHIPS and INCL constantly demonstrating more accurate predictions when compared to data, outperforming FTFP. To ensure an authentic comparison with data, the analysis of the simulated data encompassed all events, irrespective of whether antiprotons annihilated in the foil or in its mechanical support (7% of the annihilations), as previously mentioned in Sec. 1. For illustration, excluding antiprotons annihilating on the support structure from the analysis in the Monte Carlo data leads to a marginal change of about 0.2% of the average MIP multiplicity.
A comparison between measurements and simulations of the average MIPs multiplicities detected with the hodoscope are given in Table 1, along with the statistical errors. Fig. 8 illustrates these multiplicities as a function of the target atomic mass. The predicted dependence on A is slowly decreasing in FLUKA but is constant for FTFP and CHIPS. INCL constantly underestimates the MIP multiplicities while showing the strongest decrease with A among the models. Previous experimental studies for annihilation of stopped antiprotons indicate a rapid decrease in charged pion multiplicity from A=2 to A80, following an exponential law, after which it remains nearly constant for higher mass numbers [15]. In our results, the contribution from particles other than charged pions to the 3D tracks leads rather to a slowly decreasing dependence on A, in line with the FLUKA prediction. However, the disagreement between data and predictions increases with A. For FLUKA, it ranges from 1 for carbon, for molybdenum to 8 for gold. The CHIPS model predictions are almost as precise as FLUKA for carbon and molybdenum, but are off by for gold. The results obtained with the FTFP model show a discrepancy between about 7 and 20 with our experimental data. The new INCL model underestimates the MIP multiplicity between 7% and 10%.
C | Mo | Au | |
Data | 2.775(7) | 2.642(3) | 2.491(4) |
FLUKA | 2.780(6) | 2.739(6) | 2.686(6) |
FTFP | 2.981(6) | 2.973(6) | 2.987(6) |
CHIPS | 2.793(6) | 2.732(6) | 2.804(6) |
INCL | 2.580(6) | 2.356(6) | 2.306(6) |
@@
In Fig. LABEL:fig:HIPs_distros the multiplicity distributions of HIPs are compared between data and Monte Carlo simulations for the three target foils. Each model is most successful in describing the HIP multiplicity for carbon, the lightest of the three nuclei in this work. FLUKA stands out as the model that provides the best description of the HIPs multiplicity across all three targets, followed by CHIPS. In contrast, FTFP exhibits generally poor agreement with the measured data, and this disparity intensifies with increasing atomic mass.
C | Mo | Au | |
Data | 0.521(3) | 0.730(5) | 0.598(4) |
FLUKA | 0.575(3) | 0.88(1) | 0.701(4) |
FTFP | 0.412(3) | 0.120(2) | 0.047(1) |
CHIPS | 0.679(3) | 1.290(7) | 0.268(3) |
INCL | 0.383(8) | 0.399(8) | 0.225(3) |
@@
Fig. LABEL:fig:sims_HIPs_detail provides a breakdown of the various HIPs generated in annihilation and detected by the Timepix3, focusing on FLUKA and CHIPS, the two models that most accurately describe the average multiplicity. Protons are dominant nuclear fragments in both models. In FLUKA, about 50% of HIPs for -C and -Au annihilation are protons, increasing to around 60% for -Mo. In CHIPS, protons account for nearly 80% of all HIPs in -C annihilation and exceed 95% for Mo and Au. While both models feature fragments with Z=2 as the second most represented heavy prongs, they differ by larger factors for Z2. In FLUKA, Z2 fragments constitute between 2% and 6% of the total HIPs, while in CHIPS they are similar for -C annihilation (2.5%), but drop to 0.1% for molybdenum and gold.
@@ 4.2 Energy deposit from heavy prongs
The energy deposit from a HIP inside the silicon sensor of the Timepix3 quad detector depends on the type of particle, its kinetic energy as well as the angle under which the particle hits the detector (Sec. 2.1). The comparison between data and Monte Carlo models encompasses all HIPs without distinction, and is shown in Fig. LABEL:fig:HIPs_Etot. The bottom part of the plots shows the difference between data and models in terms of the standard deviation units. As detailed in Sec. 3, the volcano effect, an artifact of the Timepix3 ASIC, constrains the comparison up to 5 MeV.
@@ 5 Conclusions
To validate Monte Carlo simulation models in the widely-used packages Geant4 and FLUKA, we measured charged particles emitted from the annihilation of antiprotons in carbon, molybdenum, and gold targets. The comparison showed that overall, FLUKA demonstrated the best performance in describing the production of MIPs and HIPs, two quantities tightly bound to the final state interactions between initially produced mesons and residual nuclei. FLUKA’s predictions agree with experimental data within about 20% relative precision, which was confirmed also in ref. [44]. Among the Geant4 models, CHIPS and INCL show similar performance, outperforming FTFP in reproducing fragment multiplicities. However, CHIPS has been discontinued in the latest versions of Geant4 and is no longer maintained. The primary drawback of FTFP is the underestimation of HIP production by large factors, ranging up to 12. Despite the low rate of produced HIPs, its predictions of the deposited energy of heavy prongs are the most accurate among the examined models.
@@
Annihilation data at rest using new technologies are needed to test the annihilation mechanisms [82]. Future measurements aiming to study the final state interactions will require a more detailed identification of the types of outgoing heavily ionizing particles, for which limited data currently exists. Measuring annihilations in nuclei containing few nucleons, such as Be or C with this approach can efficiently identify particular FSIs of the annihilation mesons and nucleons. This will be considered in our next project, involving a detailed, systematic study of A annihilation at rest, covering approximately 15 different nuclei. It will be conducted at the AD/ELENA [83] facility at CERN, using the newly developed, dedicated beamline for slow extracted antiprotons at the ASACUSA experiment. The total multiplicity in nearly 4 solid angle for charged prongs of various types will be measured, along with their kinetic energy and angular distribution. The project will provide an in-depth investigation of the FSIs, which evolve with the atomic number, enhancing our understanding of antiproton-nucleus interactions at low energies. Moreover, the acquired dataset will be used for detailed validation of the current, but also future models describing antiproton-nucleus annihilation at rest.
Acknowledgments
@@
This work is supported by JSPS KAKENHI Fostering Joint International Research No. B 19KK0075, No. A 20KK0305, Grant-in-Aid for Scientific Research No. B 20H01930, No. A 20H00150; the Austrian Science Fund (FWF) Grant Nos. P 32468, W1252-N27, and P 34438; Special Research Projects for Basic Science of RIKEN; Università di Brescia and Istituto Nazionale di Fisica Nucleare; the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie Grant Agreement No. 721559; the Research Grants Program of the Royal Society and the Foundation for Dutch Scientific Research Institutes. The authors express their gratitude to the Medipix Collaboration at CERN for providing the Timepix3 ASICs for the measurements, and to Alberto Ribon from the Geant4 group at CERN for the fruitful discussions and advises. This research was funded in whole or in part by the Austrian Science Fund (FWF) [P 34438]. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.

### v2

@@ 3 Monte Carlo simulations
The Monte Carlo simulations were performed using three of the physics lists available in Geant4, as well as FLUKA (fluka 4-2.1). The model was simulated in Geant4.9.6.p04, in Geant4.10.05.p01, and the most recent one, in Geant4.11.2.1. The four models adopt different approaches when describing antiproton-nucleus annihilation at rest. The presented comparisons comprise the aforementioned physics lists in their integral form, but in future this work could be used to compare the relative importance of different effects.
@@
For the hodoscope a straightforward threshold on the deposited energy was implemented to define a hit (0.7 MeV for the bars and of 0.55 MeV for fibres), determined from the deposited energy by simulated cosmic rays. Both Geant4 and FLUKA simulations consistently indicate that most of the tracks in the hodoscope are from charged pions. However, a significant portion is ascribed to other high–energy particles crossing the four layers of the hodoscope. In FTFP, CHIPS and INCL their contribution remains constant at 35%, 50% and 40% respectively, irrespective of the target. In contrast, for FLUKA the values range between 25% and 35%, increasing with the mass of the nucleus. In CHIPS and FTFP models 95% of the non-pion tracks are generated by electrons or positrons, each with tens of MeV of energy, resulting from the conversion of -rays emitted from the decay of from annihilation. In FLUKA however, the surplus of 3D reconstructed tracks is primarily attributed to high-energy protons. In INCL most of the non-pion tracks originate from electrons or positron, but proton contribution is significant with 40%. Fig. 9 shows the distribution of the number of 3D tracks per annihilation event for -Au, across the four models. Additionally, the graph includes a breakdown of the particles fully traversing the hodoscope, depositing in each layer energy above the threshold that defines a hit. It is evident that the distribution of charged pions alone does not reproduce the number of 3D reconstructed tracks, underscoring the contribution of other particles crossing the detector. This becomes particularly important when discussing the multiplicity of MIPs, as it should be noted that in the annihilation process, particles other than charged pions, such as protons and electrons, can fully traverse the Hodoscope and be counted as MIPs.
@@
The Timepix3 simulated data were digitised using the modular simulation framework, applying the same bias voltage and charge threshold as during the actual measurements [75, 76]. The clusters generated from MIPs in the cosmic rays simulations reproduced the morphology and the total cluster energy of the measured ones, as observed in Fig. 10a. On the other hand, the digitisation could not consistently reproduce the measured signal from HIPs, due to features of the Timepix3 ASIC, such as the halo, plasma and the volcano effect [60, 77, 78, 79, 80]. These effects arise from substantial energy deposits, and simulating them accurately is challenging because HIPs do not always trigger each of them. The halo (or ”skirt”), indicating induced low energy signal in the neighboring pixels is consistently present. However, when heavier or more energetic ions deposit a significant amount of charge over short distances, two additional phenomena may occur: the plasma effect, which extends the charge collection time and consequently increases the track width, or the volcano effect, where pixels with energy deposits exceeding 500 keV register significantly lower values. The measured cluster energy is then notably lower than the actual deposited energy in the silicon sensor. In addition, the stochastic nature, in particular of the volcano effect, makes their qualitative and quantitative replication in the digitisation unfeasible. Therefore, these effects had to be individually addressed in the data analysis.
@@
The halo pixels, which are easily identifiable due to the very low energy deposits, were excluded implementing an energy cutoff of 5 keV, based on the pixel energy distribution derived from data. Fig. 10b shows a good agreement of the energy deposition by particles from a source between data (after halo removal) and Geant4 simulation, both before and after digitisation, despite the significantly larger clusters observed in data due to the presence of halo pixels. The small discrepancy is likely attributed to the uncertainties in measuring the distance between the source and the detector during the measurements.
@@ 4.1 Multiplicities
The charged pion multiplicity in antiproton-nucleus annihilation at rest is influenced by the final state interactions of the primary pions with the remaining nucleus. The number of pions that interact with the nucleus is predominantly determined by simple geometric factors, such as the proximity of the annihilation point to the nuclear surface and the size (mass) of the nucleus. The number of surviving, and detectable pions which are not absorbed by the nucleus is related to the primary pions through:
@@
Due to the limited geometrical acceptance and efficiency of the hodoscope, only a fraction of the surviving charged pions are measured in this experiment. The distributions of MIP particles presented through the number of reconstructed 3D tracks in the hodoscope are shown in Fig. 12. No model reproduces accurately the distributions obtained from the measurements for all three targets, with FLUKA, CHIPS and INCL constantly demonstrating more accurate predictions when compared to data, outperforming FTFP. To ensure an authentic comparison with data, the analysis of the simulated data encompassed all events, irrespective of whether antiprotons annihilated in the foil or in its mechanical support (7% of the annihilations), as previously mentioned in Sec. 1. For illustration, excluding antiprotons annihilating on the support structure from the analysis in the Monte Carlo data leads to a marginal change of about 0.2% of the average MIP multiplicity.
A comparison between measurements and simulations of the average MIPs multiplicities detected with the hodoscope are given in Table 1, along with the statistical uncertainties. Fig. 13 illustrates these multiplicities as a function of the target atomic mass. The predicted dependence on A is slowly decreasing in FLUKA but is constant for FTFP and CHIPS. INCL constantly underestimates the MIP multiplicities while showing the strongest decrease with A among the models. Previous experimental studies for annihilation of stopped antiprotons indicate a rapid decrease in charged pion multiplicity from A=2 to A80, followed by exponential law, after which it remains nearly constant for higher mass numbers [15]. In our results, the contribution from particles other than charged pions to the 3D tracks leads rather to a slowly decreasing dependence on A, in line with the FLUKA prediction. However, the disagreement between data and predictions increases with A. For FLUKA, it ranges from 1 for carbon, for molybdenum to 8 for gold. The CHIPS model predictions are almost as precise as FLUKA for carbon and molybdenum, but are off by for gold. The results obtained with the FTFP model show a discrepancy between about 7 and 20 with our experimental data. The new INCL model underestimates the MIP multiplicity between 7% and 10%.
[…]
@@
In Fig. 14 the multiplicity distributions of HIPs are compared between data and Monte Carlo simulations for the three target foils. Each model is most successful in describing the HIP multiplicity for carbon, the lightest of the three nuclei in this work. FLUKA stands out as the model that provides the best description of the HIPs multiplicity across all three targets, followed by CHIPS. In contrast, FTFP exhibits generally poor agreement with the measured data, and this disparity intensifies with increasing atomic mass.
[…]
@@
Fig. 15 provides a breakdown of the various HIPs generated in annihilation and detected by the Timepix3, focusing on FLUKA and CHIPS, the two models that most accurately describe the average multiplicity. Protons are dominant nuclear fragments in both models. In FLUKA, about 50% of HIPs for -C and -Au annihilation are protons, increasing to around 60% for -Mo. In CHIPS, protons account for nearly 80% of all HIPs in -C annihilation and exceed 95% for Mo and Au. While both models feature fragments with Z=2 as the second most represented heavy prongs, they differ by larger factors for Z2. In FLUKA, Z2 fragments constitute between 2% and 6% of the total HIPs, while in CHIPS they are similar for -C annihilation (2.5%), but drop to 0.1% for molybdenum and gold.
@@ 4.2 Energy deposit from heavy prongs
The energy deposit from a HIP inside the silicon sensor of the Timepix3 quad detector depends on the type of particle, its kinetic energy as well as the angle under which the particle hits the detector (Sec. 2.1). The comparison between data and Monte Carlo models encompasses all HIPs without distinction, and is shown in Fig. 16. The bottom part of the plots shows the difference between data and models in terms of the standard deviation units. As detailed in Sec. 3, the volcano effect, an artifact of the Timepix3 ASIC, constrains the comparison up to 5 MeV.
@@ 5 Conclusions
To validate Monte Carlo simulation models in the widely-used packages Geant4 and FLUKA, we measured charged particles emitted from the annihilation of antiprotons in carbon, molybdenum, and gold targets. The comparison showed that overall, FLUKA demonstrated the best performance in describing the production of MIPs and HIPs, two quantities tightly bound to the final state interactions between initially produced mesons and residual nuclei. FLUKA’s predictions agree with experimental data within about 20% relative precision, which was confirmed also in ref. [45]. The discrepancy between the measured multiplicity of HIPs and FLUKA for Au in our study (15%) is larger than what was reported in their study (where they were in agreement). A possible explanation for this difference is the variation in target thickness–1 in our case versus 10 in theirs. The thinner target allows to detect HIPs over a broader energy range, including protons with energies 2 MeV and alpha particles between 2 MeV and 10 MeV, which cannot leave the 10 target. For heavier ions, this energy range is even wider; for example, for carbon ions, it extends from 5 MeV to 50 MeV. Therefore, the differing levels of agreement between FLUKA and experimental data are likely due to FLUKA overestimating their production.
Among the Geant4 models, CHIPS and INCL show similar performance, outperforming FTFP in reproducing fragment multiplicities. However, CHIPS has been discontinued in the latest versions of Geant4 and is no longer maintained. The primary drawback of FTFP is the underestimation of HIP production by large factors, ranging up to 12. Despite the low rate of produced HIPs, its predictions of the deposited energy of heavy prongs are the most accurate among the examined models.
@@
Annihilation data at rest using new technologies are needed to test the annihilation mechanisms [84]. Future measurements aiming to study the final state interactions will require a more detailed identification of the types of outgoing heavily ionizing particles, for which limited data currently exists. Measuring annihilations in nuclei containing few nucleons, such as Be or C, with this approach can efficiently identify particular FSIs of the annihilation mesons and nucleons. This will be considered in our next project, involving a detailed, systematic study of A annihilation at rest, covering approximately 15 different nuclei. It will be conducted at the AD/ELENA facility at CERN, using the newly developed, dedicated beamline for slow extracted antiprotons at the ASACUSA experiment. The total multiplicity in nearly 4 solid angle for charged prongs of various types will be measured, along with their kinetic energy and angular distribution. The project will provide an in-depth investigation of the FSIs, which evolve with the atomic number, enhancing our understanding of antiproton-nucleus interactions at low energies. Moreover, the acquired dataset will be used for detailed validation of the current, but also future models describing antiproton-nucleus annihilation at rest.
Acknowledgements
@@
This work is supported by JSPS KAKENHI Fostering Joint International Research No. B 19KK0075, No. A 20KK0305, Grant-in-Aid for Scientific Research No. B 20H01930, No. A 20H00150; the Austrian Science Fund (FWF) Grant Nos. P 32468, W1252-N27; Special Research Projects for Basic Science of RIKEN; Università di Brescia and Istituto Nazionale di Fisica Nucleare; the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie Grant Agreement No. 721559; the Research Grants Program of the Royal Society and the Foundation for Dutch Scientific Research Institutes. The authors express their gratitude to the Medipix Collaboration at CERN for providing the Timepix3 ASICs for the measurements, and to Alberto Ribon from the Geant4 group at CERN for the fruitful discussions and advises.
This research was funded in whole or in part by the Austrian Science Fund (FWF) P 34438. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.